# Active Preference Learning for Ordering Items In- and Out-of-sample

### v1

@@ 4 A bound on the ordering error
A proof is given in Appendix D.
Theorem 1 suggests that the probability of $R(\theta_{T})\geq\epsilon$ decays exponentially with a rate that depends on the quantities $\max_{i,j}\dot{\sigma}(z_{ij}^{\top}\theta_{T})\|z_{ij}\|_{\tilde{\mathbf{H}}^{-1}_{t}(\theta_{*})}$ and $\max_{i,j}\|z_{ij}\|^{2}_{\tilde{\mathbf{H}}^{-1}_{t}(\theta_{*})}$. Both quantities are random variables that depend on the particular sampling strategy that yields $\mathbf{H}_{t}$. Focusing on the leading term, $\max_{i,j}\dot{\sigma}(z_{ij}^{\top}\theta_{T})\|z_{ij}\|_{\tilde{\mathbf{H}}^{-1}_{t}(\theta_{*})}$, Theorem 1 suggests that an active learner should gather data to minimize this quantity and obtain the smallest possible bound.
The factor $\|z_{ij}\|^{2}_{\tilde{\mathbf{H}}^{-1}_{T}(\theta_{T})}$, is related to active sampling criteria based on Fisher information (Kirsch and Gal, 2022); it is the weighted norm of $z_{ij}$ w.r.t. the inverse of the observed Fisher information. It controls the shape of the confidence ellipsoid around $\theta_{T}$ and the width of the confidence interval around $\theta_{T}^{\top}z_{ij}$. The leading term in Theorem 1 re-scales this quantity with aleatoric noise under the MLE estimate $\theta_{T}$.
Regularized maximum-likelihood.
Throughout this section, we have assumed that $\theta_{T}$ is the maximum likelihood estimate and that $\mathbf{H}(\theta_{T})$ has full rank. This can be relaxed by considering $\ell_{2}$ (Ridge) regularization where $\theta_{\lambda_{0},T}$ is the optimum of the regularized log-likelihood with regularization $\lambda_{0}\mathbf{I}$ and $\mathbf{H}_{\lambda_{0}}(\theta_{\lambda_{0},T})=\sum_{s=1}^{T}\dot{\sigma}(z_{s}^{\top}\theta_{\lambda_{0},T})z_{s}z_{s}^{\top}+\lambda_{0}\mathbf{I}$. The same machinery used to prove Lemma 1 (Filippi et al., 2010; Faury et al., 2020) can be applied to this regularized version and result in small changes to the final bound.
Other link functions.
We have focused on the logistic regression model as it is often used to model preferences. It is possible to derive similar results for other generalized linear models, with other link functions, $\mu(z_{ij}^{\top}\theta_{*})$. For this, we could use the inequality $\mathbf{H}(\theta)\geq\kappa^{-1}\mathbf{V}$ with $\mathbf{V}=\sum_{s=1}^{T}z_{s}z_{s}^{\top}$ and $\kappa\geq 1/\min_{z_{ij}}\dot{\mu}(z_{ij}^{\top}\theta_{*})$. To the best of our knowledge, this would yield a scaling of $\sim\exp(-\Delta^{2}T/\kappa)$ where, unfortunately, $\kappa$ might be very large. For discussion on the dependence on $\kappa$ in generalized linear bandits, see Lattimore and Szepesvári (2020, Chapter 19).
5 Algorithms
We present two active preference learning algorithms for ordering a list of items: GURO, a deterministic algorithm based on greedy minimization of the bound in Theorem 1, and BayesGURO, a stochastic version where we sample $\theta$ from a posterior distribution. Both are presented in Algorithm 1. As we will see, they both lead to similar sampling criteria. We begin by studying preference models in the form $\sigma(\theta^{\top}z_{ij})$ and return in Section 5.3 to variations on the model parameterization.
Algorithm 1 Greedy Uncertainty Reduction for Ordering
@@ 5 Algorithms
4:     Draw $(i_{t},j_{t})$ according to (4) or (5)
@@
7:     if GURO then
8:          Let $\theta_{t}=\mathrm{MLE}(D_{t})$,    (9) in the Appendix
9:     else if BayesGURO then
10:          Let $\theta_{t}=\mathrm{MAP}(D_{t})$,    (10) in the Appendix
11:     end if
12:end for
13:Return $h_{T}$
5.1 Greedy worst-pair uncertainty reduction
The main component of the bound in Theorem 1 to be controlled by an active learning algorithm is
@@ 5.1 Greedy worst-pair uncertainty reduction
which represents the highest uncertainty in the comparison of any items $i,j\in\mathcal{I}$ under the model $\theta_{T}$. A smaller value of (3) yields a smaller bound and a stronger guarantee.
Recall that, for any $t=1,...,T$, $\theta_{t}$ is the MLE estimate of the ground-truth parameter $\theta_{*}$ with respect to the observed history $D_{t}$. Both factors in (3) are determined by the sampling strategy that yielded the item pairs $(i_{t},j_{t})$ in $D_{T}$ and, therefore, $\mathbf{H}_{T}$ and $\theta_{T}$ (the results of comparisons $c_{ij}$ are outside the control of the algorithm, but $z_{ij}$ are known).
Direct minimization of (3) is not feasible without access to comparisons $c_{ij}$ and their likelihood under $\theta_{T}$. Instead, we adopt a greedy, alternating approach: In each round, a) a single pair is sampled for comparison by maximizing (3) under the current model estimate, and b) $\theta_{t}$ is recomputed based on $D_{t}$. Specifically, at $t=1,...,T$, we sample,
@@
We refer to this criterion as GURO for Greedy Uncertainty Reduction for Ordering, since sampling in this way reduces the uncertainty of $\theta_{t}$ in the direction of $z_{ij}$. To see this, consider the change of $\mathbf{H}_{t}(\theta_{t})$ after a single play of $i_{t},j_{t}$. The Sherman-Morrison formula (Sherman and Morrison, 1950) yields, with $z_{t}\coloneqq z_{i_{t}j_{t}}$,
$\mathbf{H}_{t}^{-1}(\theta_{t-1})$ $=\mathbf{H}_{t-1}^{-1}(\theta_{t-1})-\underbrace{\dot{\sigma}(z_{t}^{\top}\theta_{t-1})\frac{\mathbf{H}_{t-1}^{-1}(\theta_{t-1})z_{t}z_{t}^{\top}\mathbf{H}_{t-1}^{-1}(\theta_{t-1})}{1+\dot{\sigma}(z_{t}^{\top}\theta_{t-1})\|z_{t}\|^{2}_{\mathbf{H}_{t-1}^{-1}(\theta_{t-1})}}}_{=:{\color[rgb]{1,0,0}\xi}}$
@@
and thus, for all $i<j\in\mathcal{I}$, with $\mathbf{H}_{t-1}=\mathbf{H}_{t-1}(\theta_{t-1})$, it holds that $||z_{ij}||^{2}_{\mathbf{H}_{t}^{-1}(\theta_{t-1})}=||z_{ij}||^{2}_{\mathbf{H}_{t-1}^{-1}}-||z_{ij}||^{2}_{{\color[rgb]{1,0,0}\xi}}\leq||z_{ij}||^{2}_{\mathbf{H}_{t-1}^{-1}}~{}.$ The inequality is strict for the pair $i_{t},j_{t}$ in (4). As $\theta_{t}$ converges to $\theta_{*}$, this pair becomes representative of the maximizer of (3) provided there is no major systematic discrepancy between $\mathcal{I}_{D}$ and $\mathcal{I}$.
A drawback of a deterministic greedy criterion like (3) is that it cannot readily be applied in batched settings since it will sample the same pair of items throughout the batch. Next, we offer a Bayesian stochastic alternative.
5.2 Stochastic Bayesian uncertainty reduction
Consider a Bayesian model of the parameter $\theta$ with $p(\theta)$ the prior belief and $p(\theta\mid D_{t})$ the posterior after observing the preference feedback in $D_{t}$. In our second algorithm, we sample items $i_{t},j_{t}$ based on a stochastic estimate of the posterior variance of predictions for their comparison,
Here, $\hat{\mathbb{V}}_{\theta\mid D_{t}}[\sigma(\theta^{T}z_{ij})]$ is an estimate of the variance of probabilities $\sigma(\theta^{T}z_{ij})$, computed from finite samples drawn from the posterior of $\theta$. Estimating the variance in this way both i) allows for tractable implementation, and ii) induces controlled stochasticity in the selection of item pairs. We refer to the sampling criterion in (5) as BayesGURO.
For the logistic model considered in this section, using Laplace approximation with a Normal prior $\mathcal{N}(0,\mathbf{H}^{-1}_{B,0})$ on $\theta$, the Bayesian criterion in (5) is related to the criterion in (4) through the first-order Taylor expansion of the variance:
$\mathbb{V}_{\theta\mid D_{t}}(\sigma(\theta^{\top}z_{ij}))$ $\approx(\dot{\sigma}(\mathbb{E}_{\theta\mid D_{t}}[\theta^{\top}z_{ij}]))^{2}\mathbb{V}_{\theta\mid D_{t}}[\theta^{\top}z_{ij}]$
$=(\dot{\sigma}(\theta_{B,t}^{\top}z_{ij})||z_{ij}||_{\mathbf{H}^{-1}_{B,t}(\theta_{B,t})})^{2}~{},$
where $\theta_{B,t}$ is the MAP estimate of $\theta$ at time $t$ and $\mathbf{H}_{B,t}$ is the Hessian adjusted by the prior covariance $\mathbf{H}^{-1}_{B,0}$ (further described in Appendix C.2). Thus, to a first-order approximation, for a large number of posterior samples, the two active learning criteria are equivalent, save for the influence of the prior. In practice, we find that the Bayesian variant lends itself well to sequential updates of the posterior. The choice of prior $p(\theta)$, which could be useful under strong domain knowledge, and the stochasticity of using few posterior samples to approximate $\mathbb{V}$ make the two criteria distinct.
5.3 Preference models
We consider logistic preference models aimed at mitigating different sources of error. In all variants, we let $h(i,j)=\mathds{1}[f(i,j)>0]$ for some function $f$ approximating the difference in score between items $i$ and $j$. Throughout, we estimate $f$ using probabilistic modeling, with a logistic likelihood function $\sigma(f(i,j))\approx p(C_{ij}=1)$, matching (2).
@@ 5.3 Preference models
Our default model (see (2) and Section 4) is fully contextual,
$f_{\theta}(x_{i},x_{j})=\theta^{\top}(x_{i}-x_{j})~{}.$ (6)
@@
A strength of this model is that the variance in its estimates grows with $d$, but not with $n=|\mathcal{I}|$. As we see empirically (e.g., in Figure 4), this results in much quicker convergence than non-contextual methods. This fully contextual model also generalizes to unseen items in $\mathcal{I}$ under modest assumptions on how $\mathcal{I}_{D}$ was sampled.
The model’s main drawback is two other sources of error: model misspecification bias (error due to the chosen functional form), and noise (error due to insufficiency of $X$ in predicting $C$). The former can be mitigated by applying the linear model to a representation function $\phi:\mathcal{X}\rightarrow\mathbb{R}^{d^{\prime}}$,
$f_{\theta}(x_{i},x_{j})=\theta^{\top}(\phi(x_{i})-\phi(x_{j}))~{}.$
@@
A good representation $\phi$, e.g., from a foundation model, can mitigate misspecification bias, and enable using different data modalities of input. As demonstrated in Figure 7 in the Appendix, even a representation pre-trained for a different task can perform much better than a random initialization. [^1] It is also feasible to update representations during exploration (Xu et al., 2022), but we do not consider that here.
Noise due to insufficiencies in $X$ cannot be mitigated by a representation $\phi(x)$; If $X$ does not sufficiently cover aspects considered by annotators, even the Bayes classifier $\mathds{1}[p(C_{ij}\mid X_{i},X_{j})>0.5]$ will have a large ordering error. However, for in-sample ordering, we can introduce per-item parameters $\zeta_{i}\in\mathbb{R}$ to the score function, one for each item $i\in\mathcal{I}_{D}$, that compensate for both model misspecification and for uninformative attributes,
@@
Here, $\zeta_{i}-\zeta_{j}$ corrects the residual of the fully contextual model, which can be assumed small if a) the context captures the most relevant information about the ordering, and b) the functional form $\theta^{\top}\phi(x_{i})$ is nearly well-specified. $\zeta_{i}-\zeta_{j}$ alone are sufficient to learn a perfect in-sample ordering, at the cost of increased variance (the full dimension is now $n$ instead of $d$) and poor generalization (since $\zeta_{i}$ are unknown for items $i\not\in\mathcal{I}_{D}$). In experiments, the hybrid model in (7) exhibits the best of both contextual and per-item parameterizations: When $\phi$ is poor, the model recovers and competes with non-contextual models (Figure 7); when $\phi$ is good, convergence is as quick as fully contextual models.
@@ 6 Experiments
We evaluate the GURO variants in two image ordering tasks, one with logistic (synthetic) preference feedback, and one based on real-world feedback from human annotators. We provide a synthetic experiment in Appendix F.3 that includes empirical estimates of the bound in Theorem 1.
Baselines
We select diverse baseline algorithms for active sampling of comparisons: BALD (Houlsby et al., 2011), CoLSTIM (Bengs et al., 2022), TrueSkill (Graepel, 2012; Hees et al., 2016), Uniform sampling, and a version of GURO that ignores aleatoric uncertainty (referred to as NormMin).
Bayesian Active Learning by Disagreements (BALD) is, a priori, considered the strongest baseline since it is a contextual active learning algorithm, unlike the others. When applied to preference learning, BALD determines the next comparison by maximizing the following criterion
$\operatorname*{arg\,max}_{i\neq j\in\mathcal{I}_{D}}H[C_{ij}\mid z_{ij},D_{t}]-\underset{p(\theta\mid D_{t})}{\mathbb{E}}[H[C_{ij}\mid z_{ij},\theta_{t}]]~{},$ (8)
where $H$ is the conditional entropy. The criterion aims to greedily maximize the decrease in posterior entropy, which amounts to reducing the epistemic uncertainty. The second term in (8) will downplay the influence of the aleatoric uncertainty. This is not always beneficial, as suggested by our results in Section 4, since learners may require several comparisons of high-uncertainty pairs to get the order right.
CoLSTIM was developed for regret minimization and is not expected to perform well here, but is included to illustrate the mismatch between regret minimization and our setting. Uniform sampling is oblivious to the observed data and represents a non-active learner, but remains a competitive baseline in active learning tasks (Yang and Loog, 2018).
@@
TrueSkill (Herbrich et al., 2006) is a non-contextual skill-rating system that models each item as a Gaussian distribution without leveraging item attributes. We use the sampling rule from Hees et al. (2016), designed for ordering. Finally, to illustrate the importance of accounting for aleatoric uncertainty, we use a version of GURO called NormMin that ignores the $\dot{\sigma}(z_{ij}^{\top}\theta_{t})$ term and plays the pair maximizing $\|z_{ij}\|_{\mathbf{H}_{t}^{-1}(\theta_{t})}$, i.e., it minimizes the second-order term in Lemma 1. See Appendix F.3 for a detailed comparison.
(a) Kendall’s tau distance between estimated and true orderings.
(b) Mean generalization error with the 95% confidence interval.
Figure 1: X-RayAge. Performance of active sampling strategies when comparisons are simulated using a logistic regression model according to (2). In-sample Kendall’s Tau distance $R_{I_{D}}$ on 200 images (left) and generalization error $R_{I_{E}}-R_{I_{D}}$ for models trained on 150 images and evaluated on 150 images from a different distribution (right). All results are averaged over $100$ different random seeds.
Implementation details
GURO, CoLSTIM, and Uniform use the logistic regression implementation in Scikit-learn (Pedregosa et al., 2011) with default Ridge regularization ($C=1$). The former two update $\theta_{t}$ every iteration using the full history, $D_{t}$. The design matrix for CoLSTIM is updated as in Bengs et al. (2022): the confidence width $c_{1}$ was chosen to be $\sqrt{d\log(T)}$, and the perturbed values were generated using the standard Gumbel distribution. GURO updates $\mathbf{H}^{-1}_{t}$ every iteration using the Sherman-Morrison formula.
For BayesGURO and BALD, the posterior $p(\theta\mid D_{t})$ is estimated using the Laplace approximation as described in Bishop and Nasrabadi (2006). With this approximation, the covariance matrix is the same as the inverse of the Hessian of the log-likelihood. For both methods, the priors $\theta_{B,0}=\mathbf{0}^{d}$ and $\mathbf{H}_{B,0}^{-1}=I_{d}$ were used, and sequential updates were performed every iteration. The sample criterion for BALD under a logistic model is given in Appendix C.4.1. For BayesGURO, 50 posterior samples were used to estimate $\hat{\mathbb{V}}_{\theta\mid D_{t}}[\sigma(\theta^{T}z_{ij})]$ for every $z_{ij}$. The hybrid algorithms follow the same structure, with a small change in sampling (see Appendix F.2), and a constraint that each per-item parameter $\zeta_{i}$ is independent of other parameters. This allows for efficient updates of $\mathbf{H}_{B,t}^{-1}$ by using sparsity in the covariance.

### v2

@@ 4 Which comparisons result in a good ordering?
Theorem 1 suggests that the probability of $R(\theta_{T})\geq\epsilon$ decays exponentially with a rate that depends on the quantities $\max_{i,j}\dot{\sigma}(z_{ij}^{\top}\theta_{T})\|z_{ij}\|_{\tilde{\mathbf{H}}^{-1}_{t}(\theta_{T})}$ and $\max_{i,j}\|z_{ij}\|^{2}_{\tilde{\mathbf{H}}^{-1}_{t}(\theta_{T})}$. Both quantities are random variables that depend on the particular sampling strategy that yields $\mathbf{H}_{t}$. Focusing on the leading term, $\max_{i,j}\dot{\sigma}(z_{ij}^{\top}\theta_{T})\|z_{ij}\|_{\tilde{\mathbf{H}}^{-1}_{t}(\theta_{T})}$, Theorem 1 suggests that an active learner should gather data to minimize this quantity and obtain the smallest possible bound. The factor $\|z_{ij}\|^{2}_{\tilde{\mathbf{H}}^{-1}_{T}(\theta_{T})}$ is the weighted norm of $z_{ij}$ w.r.t. the inverse of the observed Fisher information (cf. Kirsch and Gal (2022)). It controls the shape of the confidence ellipsoid around $\theta_{T}$ and the width of the confidence interval around $\theta_{T}^{\top}z_{ij}$. The leading term in Theorem 1 re-scales this quantity with aleatoric noise under the MLE estimate $\theta_{T}$. This suggests that higher epistemic (model) certainty is needed in directions with high aleatoric uncertainty—where item similarity increases noise in comparisons.
In Appendix C.3, we comment on i) generalizations to regularized preference models, ii) applications to generalized linear models with other link functions, iii) lower bounds on the ordering error, and iv) an algorithm-specific upper bound.
5 Greedy uncertainty reduction for ordering (GURO)
We present an active preference learning algorithm based on greedy minimization of the bound in Theorem 1, called GURO. We begin with fully contextual preference models of the form $\sigma(\theta^{\top}z_{ij})$ and return in Section 5.1 to parameterization variants to reduce the effects of model misspecification.
Algorithm 1 Greedy Uncertainty Reduction for Ordering (GURO), [BayesGURO]
@@ 5 Greedy uncertainty reduction for ordering (GURO)
4:     Draw $(i_{t},j_{t})$ based on $\theta_{t}$ according to (5) [or (7)]
@@
7:     $\theta_{t}=\mathrm{MLE}(D_{t})$ according to (3) [or $\theta_{t}=\mathrm{MAP}(D_{t})$ as in (11) in the Appendix]
8:end for
9:Return $h_{T}$
The main component of the bound in Theorem 1 to be controlled by an active learner is the term
@@
which represents the highest uncertainty in the comparison of any items $i,j\in\mathcal{I}$ under the model $\theta_{T}$. A smaller value of (4) yields a smaller bound and a stronger guarantee. Recall that, for any $t=1,...,T$, $\theta_{t}$ is the MLE estimate of the ground-truth parameter $\theta_{*}$ with respect to the observed history $D_{t}$. Both factors in (4) are determined by the sampling strategy that yielded the item pairs $(i_{t},j_{t})$ in $D_{T}$ and, therefore, $\mathbf{H}_{T}$ and $\theta_{T}$ (the results of comparisons $c_{ij}$ are outside the control of the algorithm, but $z_{ij}$ are known).
Direct minimization of (4), for a subset $\mathcal{I}_{D}$, is not feasible without access to comparisons $c_{ij}$ and their likelihood under $\theta_{T}$. Instead, we adopt a greedy, alternating approach: In each round, a) a single pair is sampled for comparison by maximizing (4) under the current model estimate, and b) $\theta_{t}$ is recomputed based on $D_{t}$. Specifically, at $t=1,...,T$, we sample,
@@
We refer to this sampling criterion as Greedy Uncertainty Reduction for Ordering (GURO), since it reduces the uncertainty of $\theta_{t}$ in the direction of $z_{ij}$. To see this, consider the change of $\mathbf{H}_{t}(\theta_{t})$ after a single play of $i_{t},j_{t}$. The Sherman-Morrison formula (Sherman and Morrison, 1950) yields,
$\mathbf{H}_{t}^{-1}(\theta_{t-1})=\mathbf{H}_{t-1}^{-1}(\theta_{t-1})-\dot{\sigma}(z_{t}^{\top}\theta_{t-1})\frac{\mathbf{H}_{t-1}^{-1}(\theta_{t-1})z_{t}z_{t}^{\top}\mathbf{H}_{t-1}^{-1}(\theta_{t-1})}{1+\dot{\sigma}(z_{t}^{\top}\theta_{t-1})\|z_{t}\|^{2}_{\mathbf{H}_{t-1}^{-1}(\theta_{t-1})}}~{},$ (6)
@@
where $z_{t}\coloneqq z_{i_{t}j_{t}}$. With $\xi$ as the second term in (6), it holds for all $i<j\in\mathcal{I}$, with $\mathbf{H}_{t-1}=\mathbf{H}_{t-1}(\theta_{t-1})$, that $||z_{ij}||^{2}_{\mathbf{H}_{t}^{-1}(\theta_{t-1})}=||z_{ij}||^{2}_{\mathbf{H}_{t-1}^{-1}}-||z_{ij}||^{2}_{\xi}\leq||z_{ij}||^{2}_{\mathbf{H}_{t-1}^{-1}}~{}.$ The inequality is strict for the pair $i_{t},j_{t}$ in (5). As $\theta_{t}$ converges to $\theta_{*}$, this pair becomes representative of the maximizer of (4) provided there is no major systematic discrepancy between $\mathcal{I}_{D}$ and $\mathcal{I}$.
Surprisingly, GURO can also be justified from a Bayesian analysis. Consider a Bayesian model of the parameter $\theta$ with $p(\theta)$ the prior belief and $p(\theta\mid D_{t})$ the posterior after observing the preference feedback in $D_{t}$. A natural active learning strategy is to sample items $i_{t},j_{t}$ for which the model preference is highly uncertain under the posterior distribution,
$i_{t},j_{t}=\operatorname*{arg\,max}_{i,j\in\mathcal{I}_{D},i<j}\hat{\mathbb{V}}_{\theta\mid D_{t-1}}[\sigma(\theta^{\top}z_{ij})]~{},$ (7)
where $\hat{\mathbb{V}}_{\theta\mid D_{t-1}}[\sigma(\theta^{\top}z_{ij})]$ is a finite-sample estimate of the variance in predictions, computed by sampling from the posterior. In Appendix B.3, we show that the first-order Taylor expansion of the true variance is equal to the GURO criterion. Hence, we refer to sampling according to (7) as BayesGURO. Unlike GURO, BayesGURO can incorporate prior knowledge through $p(\theta)$ and benefits from controlled stochasticity through the empirical estimate $\hat{\mathbb{V}}$, which makes it appropriate for batched algorithms—a deterministic criterion would construct batches of a single item pair. Both GURO and BayesGURO are presented in Algorithm 1.
Computational Complexity:
Running the algorithms requires $O(n^{2})$ operations each iteration to evaluate the sampling criteria (Equation 5 or 7) on all possible pairs, a problem shared by many active preference learning algorithms (Qian et al., 2015; Canal et al., 2019; Houlsby et al., 2011). A way of mitigating this computational complexity is to, at each time step, sample a fixed number of comparisons and only evaluate on these, similar to the approach taken in Canal et al. (2019). When only looking at a sample of $m\ll n^{2}$ pairs, the complexity is reduced to $O(m)$. While making $m$ too small can hurt the sample complexity, we describe in Appendix E how we implemented this sub-sampling strategy to speed up computations in one of our experiments and observed no noticeable change in performance. Lastly, we want to highlight that in many realistic scenarios, the computational burden pales in comparison to the time it takes to query an annotator.
5.1 Preference models for in- and out-of-sample ordering
@@
Our default preference model $h(i,j)=\mathds{1}[f(i,j)>0]$ is based on a fully contextual scoring function
$f_{\theta}(x_{i},x_{j})=\theta^{\top}(x_{i}-x_{j})~{},$ (8)
@@
fit with a logistic likelihood $\sigma(f(i,j))\approx p(C_{ij}=1)$. The model’s strength is that the variance in its estimates grows with $d$, but not with $n=|\mathcal{I}|$, often resulting in quicker convergence than non-contextual methods for moderate dimension $d$ (see, e.g., Figure 6). The fully contextual model also generalizes to unseen items as long as the attributes for $\mathcal{I}_{D}$ span attributes observed for $\mathcal{I}$.
@@
The limitations of a fully contextual model are model misspecification (error due to the functional form), and noise (error due to $C$ not being fully determined by $X$). The former can be mitigated by applying the linear model to a representation function $\phi:\mathcal{X}\rightarrow\mathbb{R}^{d^{\prime}}$, $f_{\theta}(x_{i},x_{j})=\theta^{\top}(\phi(x_{i})-\phi(x_{j})).$ A good representation $\phi$, e.g., from a foundation model, can mitigate misspecification and admit different input modalities. As demonstrated in Figure 9 in the Appendix, even a representation pre-trained for a different task can perform much better than a random initialization. [^1] It is feasible to update representations during exploration (Xu et al., 2022; Singh and Chakraborty, 2021), but we do not consider that here.
Noise due to insufficiencies in $X$ cannot be mitigated by a representation $\phi(x)$; If annotators consistently compare items based on features $U$ not included in $X$, no function $h(X_{i},X_{j})$ can perfectly order the items. However, for in-sample ordering of $\mathcal{I}_{D}$, adding per-item parameters $\zeta_{i}\in\mathbb{R}$ to the scoring function, one for each item $i\in\mathcal{I}_{D}$, can mitigate both misspecification and noise,
@@
We call this a hybrid model and apply it in “GURO Hybrid” and baselines in experiments. The term $\zeta_{i}-\zeta_{j}$ can correct the residual of the fully contextual model, which is small if a) the context captures the most relevant information about the ordering, and b) the functional form $\theta^{\top}\phi(x_{i})$ is nearly well-specified. Using $\zeta_{i}-\zeta_{j}$ alone is sufficient in-sample, but has high variance (the dimension is $n$ instead of $d$) and poor generalization ($\zeta_{i}$ are unknown for items $i\not\in\mathcal{I}_{D}$). In practice, we use L2 regularization to prevent the model from learning an arbitrary $\theta$ by using the full expressivity of $\zeta_{i}$ (see Appendix E for details). Empirically, our hybrid models exhibit the best of both worlds: When $\phi$ is poor, the model recovers and competes with non-contextual models (Figure 9); when $\phi$ is good, convergence matches fully contextual models (Figure 6).
@@ 6 Experiments
We evaluate GURO (Algorithm 1) and GURO Hybrid (see Section 5.1) in four image ordering tasks, one with logistic (synthetic) preference feedback, and three tasks based on real-world feedback from human annotators [^2] Our code is available at: https://github.com/Healthy-AI/GURO. We provide a synthetic experiment in Appendix E.2 that includes empirical estimates of the bound in Theorem 1. The experiments include five diverse baseline algorithms, described next. BALD (Houlsby et al., 2011) is a priori the strongest baseline since it is a contextual active learning algorithm, unlike the others. Its selection criterion greedily maximizes the decrease in posterior entropy, which amounts to reducing the epistemic uncertainty and includes a term to downplay the influence of aleatoric uncertainty. This is not always beneficial, as suggested by our analysis in Section 4, since learners may require several comparisons of high-uncertainty pairs to get the order right. CoLSTIM (Bengs et al., 2022) is a contextual bandit algorithm, developed for regret minimization and is not expected to perform well here. It is included to illustrate the mismatch between regret minimization and our setting.
@@
TrueSkill (Herbrich et al., 2006; Graepel, 2012) is a non-contextual skill-rating system that models the score of each item as a Gaussian distribution, disregarding item attributes, and has been adopted in various works to score items based on subjective pairwise comparisons (Larkin et al., 2022; Naik et al., 2014; Sartori et al., 2015). We use the sampling rule from Hees et al. (2016), designed for ordering. Finally, we include Uniform sampling, and to illustrate the importance of accounting for aleatoric uncertainty, we use a version of GURO called NormMin that ignores the $\dot{\sigma}(z_{ij}^{\top}\theta_{t})$ term and plays the pair maximizing $\|z_{ij}\|_{\mathbf{H}_{t}^{-1}(\theta_{t})}$, i.e., it minimizes the second-order term in Lemma 1. NormMin corresponds to the selection criterion in the concurrent work Das et al. (2024), adapted to our problem of finding the correct ordering. We refer the reader to Appendix E.2 for a detailed comparison where NormMin performs significantly worse than Uniform on certain problem instances, and Appendix E for details regarding the implementation and the choice of hyperparameters for GURO, BayesGURO, and baselines.